OpenStudy (anonymous):

What is the sum of the first 18 terms of the arithmetic series 11 + 15 + 19 + 23 + ⋯? A: 684 B:810 C:831 D:846

OpenStudy (amoodarya):

15=11+4 19=15+4 23=19+4 so they are arithmetic sequence a1=11 r=4 a(n)=a1+(n-1)r=11+(n-1)4=11+4n-4=4n+7 s(n)=1/2 *n(a1+an)

OpenStudy (anonymous):

Wow ok I'm stupid I just realized it has a pattern format, and just checking @amoodarya it's 810?

OpenStudy (amoodarya):

s(18)=18/2 (a1+a18 ) =9(11+11+(18-1)4) =9(22+68)=9(90)=810

OpenStudy (anonymous):

YES, thanks mate!